OpenStudy (anonymous):

X=Vot+1/2at^2 solve for Vo, then a

OpenStudy (ikram002p):

ok ? and ?

OpenStudy (anonymous):

and the question?

OpenStudy (ikram002p):

\(X= V_0t+\frac{1}{2}at^2\) \(V_0=\frac{X - \frac{1}{2}at^2}{t}\) \(a=\frac{X-V_0t}{\frac{1}{2}t^2}\)

OpenStudy (ikram002p):

np :D

OpenStudy (ikram002p):

\(\Huge{\color{green}{ت}}\)